OpenStudy (anonymous):

The length of a rectangular photograph is 1 cm. less than twice the width. If the area of the photograph is 45cm^2, find the dimensions of the rectangle.

OpenStudy (atildathehun):

What is the equation to find the area of a rectangle?

OpenStudy (anonymous):

l x w

OpenStudy (mathmale):

A = L * W. (Please reserve " x " as the representation of an algebraic quantity.) Use " * " to indicate multiplication. OK. A = L * W. How would you express the width of this photograph? How would you express the length of this photograph?

OpenStudy (anonymous):

length is 1 cm

OpenStudy (anonymous):

that is less than twice the width

OpenStudy (mathmale):

No, it's more complicated than that. We don't know the width, so we represent the width by W. "The length of a rectangular photograph is 1 cm. less than twice the width." Using the variable W, express the length of the photo algebraically.

OpenStudy (anonymous):

I'm having trouble expressing this

OpenStudy (mathmale):

If the width of the photo is W, how would you express "twice the width?"

OpenStudy (anonymous):

2w

OpenStudy (mathmale):

Good. "The length of a rectangular photograph is 1 cm. less than twice the width." Using your 2W, how would you express the length, L, of the photo, algebraically?

OpenStudy (anonymous):

1 cm

OpenStudy (anonymous):

so 1 cm times ??

OpenStudy (mathmale):

"The length of a rectangular photograph is 1 cm. less than twice the width." You must start with "2W" and then do something to it. What? What does "less than" mean?

OpenStudy (anonymous):

subtract

OpenStudy (anonymous):

2w - 1

OpenStudy (anonymous):

no sorry

OpenStudy (anonymous):

2x * 1

OpenStudy (mathmale):

That 2W - 1 is correct, actually; 2x*1 is far off. "The length of a rectangular photograph is 1 cm. less than twice the width." This says: take twice the width and then subtract 1 cm from that. Symbolically, 2W - 1 = L.

OpenStudy (anonymous):

oh no

OpenStudy (anonymous):

2w-1 * 1

OpenStudy (anonymous):

that what i meant

OpenStudy (mathmale):

the correct way to represent L is L=2W-1. No more than that. In summary: Length of rectangle: L = 2W - 1 Width of rectangle: W General formula for area of a rectangle: A = L * W. What is the area of THIS particular rectangle, expressed algebraically / symbolicallyi?

OpenStudy (anonymous):

2w^2 - w

OpenStudy (anonymous):

w(2w-1)

OpenStudy (mathmale):

Good. You are told that the actual area is 45. Thus, W(2W - 1) = 45 How would you solve this for W? Hint: Re=-arrange this into a quadratic equaton.

OpenStudy (anonymous):

oh

OpenStudy (mathmale):

Please do the work.

OpenStudy (anonymous):

SO i factor 2w^2-w-45

OpenStudy (anonymous):

i do ac method?

OpenStudy (anonymous):

?

OpenStudy (atildathehun):

The (2W - 1) is in the place of L in the equation W*L=A

OpenStudy (mathmale):

Again, the area is 45, and that 45 equals W(2W-1). Multiply out: W(2W-1). Do that now, please.

OpenStudy (mathmale):

What is W*W? What is W*1? Then, putting this together, what is W(W-1)?

OpenStudy (atildathehun):

@MathDude0000 do you remember how to do trinomials?

OpenStudy (mathmale):

@MathDude0000: I'm glad to help you. However, I do expect your full attention. I undrstand that this problem has been difficult for you, but assure you it will become easier with practice. I must now excuse myself from OpenStudy. I urge you to continue working on this problem, with Atilda.